OpenStudy (anonymous):

HELP ME PLEASE ILL MEDAL Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y A. ( 3, -1/2 ) B. ( 0 , -2 ) C. ( 0 , -11/2 ) D. ( 7, -5/2 )

OpenStudy (anonymous):

Are those two different equations or just equation put together

OpenStudy (anonymous):

Two different ones

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

The first one is 11/2

OpenStudy (anonymous):

The second one is 1

OpenStudy (anonymous):

Thats not a a b c or d choice @Thehulk49

OpenStudy (aaronandyson):

4x + 2y = 11 x - 2 = 2y Make x the subject x = -2y = 2 Plug x = -2y + 2 in 4x + 2y = 11

OpenStudy (aaronandyson):

SO,we get 4(-2y+2)+2y = 11

OpenStudy (aaronandyson):

Solve for y

OpenStudy (aaronandyson):

Hint: 4*-2 = -8 4*2 = 8

OpenStudy (aaronandyson):

@jayonna15